OpenStudy (anonymous):

medal!!!1 how do u do this @misty1212

OpenStudy (misty1212):

lol idk make another good guess!

OpenStudy (anonymous):

\[What is the value of the expression \frac{ 2 }{ 4 -3}\]

OpenStudy (anonymous):

wait

OpenStudy (misty1212):

tee hee

OpenStudy (misty1212):

\[\frac{2}{4-3}\]?

OpenStudy (anonymous):

2 ------- -3 4

OpenStudy (misty1212):

no i bet it is \[\huge \frac{2}{4^{-3}}\]

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

the exponent in the denominator i negative right?

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

that means it comes up to the numerator compute \[2\times 4^3\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

my answer choices r 128 256 1/6 1/32

OpenStudy (misty1212):

let me know when you get \(128\)